### v1

# Anisotropic Quantum Annealing vs Trit Annealing

Abstract
Quantum annealing offers a promising strategy for solving complex optimization problems by encoding the solution into the ground state of a problem Hamiltonian. While most implementations rely on spin-$1/2$ systems, we explore the performance of quantum annealing on a spin-$1$ system where the problem Hamiltonian includes a single ion anisotropy term of the form $D\sum(S^{z})^{2}$. Our results reveal that for a suitable range of the anisotropy strength $D$, the spin-$1$ annealer reaches the ground state with higher fidelity. We attribute this performance to the presence of the intermediate spin level and the tunable anisotropy, which together enable the algorithm to traverse the energy landscape through smaller, incremental steps instead of a single large spin flip. This mechanism effectively lowers barriers in the configuration space and stabilizes the evolution. These findings suggest that higher spin annealers offer intrinsic advantages for robust and flexible quantum optimization, especially for problems naturally formulated with ternary decision variables.
I Introduction
Combinatorial optimization arises across physics, computer science, and engineering, with applications from materials and circuit design to logistics, portfolio optimization, and scheduling Harrison2022max3cutcoherent; Weggemans2022solvingcorrelation. Many such problems can be cast as minimizing a cost function over discrete variables, naturally mapping to the ground state of an effective Ising-type Hamiltonian with local fields and pairwise couplings Finnila1994; Kadowaki1998; Farhi2001; Das2008; Lucas2014. Candidate solutions correspond to spin configurations, while optimal or near-optimal solutions lie at low energies. Since the configuration space grows exponentially with system size, classical methods can struggle on rugged, frustrated landscapes with many nearly degenerate local minima.
Quantum annealing (QA) is a heuristic approach to such optimization problems that leverages quantum fluctuations Kadowaki1998; Farhi2001; Albash2018Review. The protocol initializes the system in the ground state of a simple driver Hamiltonian—typically a transverse field term that does not commute with the problem Hamiltonian—and then interpolates toward the problem Hamiltonian under a time-dependent schedule. In the coherent, adiabatic limit, the system follows the instantaneous ground state provided the evolution is sufficiently slow. In practice, the runtime is constrained by the minimum spectral gap encountered along the annealing path Farhi2001; Albash2018Review; Jansen2006BoundsFT; small gaps therefore require long schedules and increase susceptibility to nonadiabatic excitations, decoherence, and control errors Bottarelli2025symmetrycd.
Most theoretical and experimental QA platforms use effective spin-$1/2$ degrees of freedom (qubits), which directly encode binary decision variables Das2008; Albash2018Review; Hauke2020; Johnson2011; DenchevPRX2016; BoixoNC2016. However, many practically relevant optimization problems are intrinsically multivalued, motivating approaches that go beyond binary encoding. Examples include ternary decisions (e.g., buy/hold/sell) and formulations with variables taking values in sets such as $\{-1,0,+1\}$. On qubit based architectures, representing multivalued logic typically requires embedding each variable into multiple qubits together with penalty terms enforcing logical constraints Weggemans2022solvingcorrelation. These embeddings increase hardware overhead, reduce effective connectivity, and complicate the energy landscape, potentially shrinking relevant gaps and amplifying sensitivity to noise and control imperfections. Higher spin systems provide a natural alternative in which multilevel quantum degrees of freedom directly encode multivalued variables.
Spin-11 systems, in particular, have three eigenstates of $S^{z}$ labeled by $m_{s}=-1,0,+1$, enabling a direct representation of ternary variables without elaborate embeddings. This has motivated interest in multilevel quantum optimization and in understanding how extra levels affect adiabatic protocols Amin2013Qudits; WangHuSandersKais2020; Weggemans2022solvingcorrelation; Bottarelli2025symmetrycd. Beyond compact encoding, spin-$1$ models also introduce control parameters absent in spin-$1/2$ systems. A key example is the single ion anisotropy term $D\sum_{i}(S_{i}^{z})^{2}$, familiar from spin-$1$ magnetism such as the Blume–Capel model Blume1966; Capel1966, which tunes the relative energetics of the $m_{s}=0$ and $m_{s}=\pm 1$ states. Since $(S^{z})^{2}$ is proportional to the identity for spin-$1/2$, this anisotropy provides a genuine additional control knob for spin-$1$ annealers and gives rise to richer behavior along the annealing path.
The remainder of this paper is organized as follows. We introduce the spin-$1$ annealing Hamiltonian and trit based simulated annealing in Sec. II. Sec. III describes our numerical protocol and performance metrics. In Sec. IV, we compare the spin-$1$ annealer with classical annealing. Sec. V summarizes our findings and outlines directions for future work. Physical interpretations based on the structure of the classical energy landscape are discussed in Appendix A.
II Theoretical Framework
We consider a spin-$1$ quantum annealer, encoding each trit in the local $S^{z}$ eigenbasis $\{|+1\rangle,|0\rangle,|-1\rangle\}$, satisfying
$S^{z}|m_{z}\rangle=m_{z}|m_{z}\rangle,\quad m_{z}\in\{-1,0,+1\}.$ (1)
In this basis, the spin operators are
$S^{z}=\begin{pmatrix}1&0&0\\[2.0pt] 0&0&0\\[2.0pt] 0&0&-1\end{pmatrix},\quad S^{x}=\frac{1}{\sqrt{2}}\begin{pmatrix}0&1&0\\[2.0pt] 1&0&1\\[2.0pt] 0&1&0\end{pmatrix}$, (2)
with $S^{\alpha},S^{\beta}]=i\epsilon_{\alpha\beta\gamma}S^{\gamma}$. The system evolves under the time-dependent Hamiltonian
$H(t)=H_{P}+H_{D}(t),$ (3)
where the problem Hamiltonian is
$H_{P}$ $=-J\sum_{\langle i,j\rangle}S_{i}^{z}S_{j}^{z}-h\sum_{i}S_{i}^{z}+D\sum_{i}(S_{i}^{z})^{2},$ (4)
and the driver Hamiltonian is
$H_{D}(t)$ $=-g(t)\sum_{i}S_{i}^{x}.$ (5)
Here $S_{i}^{\alpha}$ are spin-$1$ operators acting on site $i$, and the parameters are defined as
  • $J$ is the coupling constant that determines the interaction strength between adjacent spins,
  • $h$ is the longitudinal field, representing an external field that creates an energy preference for spin alignment in the $z$ direction,
  • $D$ is the single ion anisotropy parameter,
  • $g(t)$ is the annealing schedule controlling the strength of the transverse field.
By construction, $H_{P}$ is the quantum embedding of the classical cost function: on computational basis states $|s_{1},\dots,s_{N}\rangle$ with $S_{i}^{z}|s_{i}\rangle=s_{i}|s_{i}\rangle$ and $s_{i}\in\{-1,0,+1\}$, one has
$H_{P}\,|s_{1},\dots,s_{N}\rangle=H(s)\,|s_{1},\dots,s_{N}\rangle.$ (6)
[…]
$1-P_{gs}\lesssim\max_{t}\frac{\big|\langle\psi_{1}(t)|\dot{H}(t)|\psi_{0}(t)\rangle\big|^{2}}{\Delta(t)^{4}},$ (7)
with gap $\Delta(t)=E_{1}(t)-E_{0}(t)$. Near an avoided crossing with minimum gap $\Delta_{\min}$, a two level linearization yields the Landau–Zener formula Landau
$P_{\rm exc}\approx\exp\left(-\frac{\pi\Delta_{\min}^{2}}{2\hbar|\alpha|}\right),$ (8)
where $\alpha$ is the diabatic slope difference.
The corresponding classical configuration space is
$\mathcal{S}=\{-1,0,+1\}^{N},$ (9)
with cost function
$H(s)$ $=-J\sum_{\langle i,j\rangle}s_{i}s_{j}-h\sum_{i=1}^{N}s_{i}+D\sum_{i=1}^{N}s_{i}^{2}.$ (10)
At temperature $T>0$, a single-site Metropolis update Metropolis1953 proceeds as follows:
  1. 1.
    Pick a site $i\in\{1,\dots,N\}$ uniformly at random.
  2. 2.
    Propose a new local value $s_{i}^{\prime}\in\{-1,0,+1\}\setminus\{s_{i}\}$ uniformly over the two alternatives (the proposal distribution is symmetric).
  3. 3.
    Let $s^{\prime}$ be $s$ with $s_{i}$ replaced by $s_{i}^{\prime}$. Define the local energy for setting $s_{i}=m$ by
    $\varepsilon_{i}(m)=-Jm\sum_{j\in\partial i}s_{j}-hm+Dm^{2},$ (11)
    where $\partial i$ denotes the neighbors of $i$. The associated energy change is
    $\Delta E$ $=H(s^{\prime})-H(s)$ (12)
    $=\varepsilon_{i}(s_{i}^{\prime})-\varepsilon_{i}(s_{i})$ (13)
    $=-\big(s_{i}^{\prime}-s_{i}\big)\left(J\sum_{j\in\partial i}s_{j}+h\right)+D\big[(s_{i}^{\prime})^{2}-s_{i}^{2}\big].$ (14)
  4. 4.
    Accept the proposal with Metropolis probability
    $P_{\rm acc}(s\to s^{\prime})=\min\left\{1,e^{-\Delta E/T}\right\};$ (15)
    otherwise keep $s$.
III Methods
We compare anisotropic quantum annealing (AQA) with trit annealing (TA) on a one dimensional nearest neighbor open chain of length $N=5$. Each site carries a spin–1 degree of freedom, giving a Hilbert space dimension $\dim\mathcal{H}=3^{N}$, which remains small enough to allow exact diagonalization of the problem Hamiltonian and numerically controlled evolution. To lift the degeneracy between globally spin flipped configurations while preserving the essential competition between exchange and single–ion anisotropy, we fix the longitudinal field to a small nonzero value $h=0.2$. This symmetry breaking field selects a unique classical ground state for each $(J,D)$ pair but is sufficiently weak not to qualitatively alter the underlying physics. Working at fixed small $N$ lets us disentangle intrinsic algorithmic behavior from finite size scaling issues, while the resulting spectrum still exhibits nontrivial features such as avoided crossings and associated diabatic effects.
Both AQA and TA are driven by annealing schedules that share the same overall shape, specified by a smooth function of a dimensionless time parameter. In AQA, the system evolves under a time dependent Hamiltonian $H(t)$ that goes from a transverse driver to classical problem Hamiltonian via a scalar control function $g(t)$ over a total evolution time $T$. In TA, we use the same functional form to define a cooling schedule, but now parameterized over a total of $S$ Monte Carlo sweeps, so that the classical dynamics experiences an analogous slowdown pattern. The transverse driver amplitude is scheduled as
$g(t)=\frac{e}{f(t)},\qquad e\in\{1,5,10,20\},$ (16)
with four representative decay profiles
$f(t)\in\Big\{\log(1{+}t),\ \sqrt{t},\ t,\ t^{2}\Big\}.$ (17)
[…]
$P_{AQA}=\bigl|\langle\psi_{g}|\psi(T)\rangle\bigr|^{2},$ (18)
[…]
TA is defined on the same set of spin configurations by taking the classical cost function to be the diagonal of the quantum problem Hamiltonian, so that classical and quantum instances are exactly aligned at the level of their energy landscapes. One sweep comprises $N$ single site updates, in which each spin is updated once on average according to the local Boltzmann weights at the current effective temperature. The inverse temperature is reduced following the same schedule shape $g(t)$ used in the quantum protocol, now applied over the course of $S$ sweeps. For each disorder instance (coupling realization) we perform $R$ independent TA runs starting from random initial spin configurations drawn uniformly from the $3^{N}$ classical states. At the end of each run $(r)$ we record the terminal energy $E_{\mathrm{final}}^{(r)}$. The exact classical ground state energy $E_{0}$ is obtained by exhaustive enumeration of all configurations. We define the TA success probability as
$P_{\mathrm{TA}}=\frac{1}{R}\sum_{r=1}^{R}\mathbb{I}\!\left[\left|E_{\mathrm{final}}^{(r)}-E_{0}\right|\leq 10^{-6}\right],$ (19)
where $\mathbb{I}[\cdot]$ denotes the indicator function. The small numerical tolerance accounts for floating point round off while effectively requiring that the algorithm reaches the true ground state. By matching instance sets, schedule shapes, and problem Hamiltonians between AQA and TA, we can directly compare their success probabilities.
IV Results
Figure 1: AQA versus TA across the coupling–anisotropy plane. Each panel displays the difference $P_{\mathrm{AQA}}-P_{\mathrm{TA}}$ over the $(J,D)$ plane for a fixed transverse field scale $e$ and annealing schedule $g(t)$, at longitudinal field $h=0.2$. Columns (left to right) correspond to schedules with late-time decay $g(t)\propto 1/\log(1{+}t)$, $1/\sqrt{t}$, $1/t$, and $1/t^{2}$, while rows (top to bottom) correspond to $e=1,5,10,20$. Red (blue) indicates parameter regions where AQA yields higher (lower) success probability than TA, and yellow markers indicate cases with $P_{\mathrm{AQA}}>0.9$. The dashed gray lines at $J=0$ and $D=0$ delineate the easy-plane ($D>0$) and easy-axis ($D<0$) sectors and assist in identifying symmetry-related structures.
Figure 1 compares the performance of AQA and TA across the $(J,D)$ plane for all transverse field scales and schedules considered. The dominant trends are governed by the driver strength $e$, the anisotropy sector (easy-plane $D>0$ versus easy-axis $D<0$), and the late-time decay of the schedule $g(t)$.
For weak driving ($e=1$, top row), TA is comparable to AQA over nearly the entire parameter range, as indicated by the predominance of blue regions across all schedules. In this regime, the transverse field contribution is insufficient, within the fixed runtime, to reliably generate transitions that improve the final success probability. The sparse red features occur in narrow parameter bands near boundaries between distinct classical ordering tendencies, consistent with cases where near degeneracies permit limited benefit from coherent mixing.
As the driver strength increases ($e=5$ and above), extended parameter regions emerge in which AQA outperforms TA. This advantage is most apparent in the easy-plane sector ($D>0$) at intermediate coupling magnitudes $|J|$, where red domains become substantial for $e=5$ and expand further for $e=10$ and $e=20$. The growth of these AQA favorable regions indicates that sufficiently strong transverse driving can more effectively exploit low energy structure associated with competing classical patterns.
The easy-axis sector ($D<0$) exhibits a more restrained AQA advantage. Across many parameters TA remains competitive, and broad AQA dominant regions occur less frequently than in the easy-plane case. This behavior is consistent with an effective restriction of relevant low energy configurations to the $\pm 1$-subspace when the 0-level is disfavored, which reduces the diversity of low energy pathways available to intermediate state dynamics (See App. A for more details).
Across all driver strengths, structured V-shaped and X-shaped patterns appear near order change boundaries. These boundary regions are frequently associated with reduced performance for both protocols, producing narrow corridors in which the relative advantage alternates and separating neighboring domains with more uniform behavior.
At fixed e, performance exhibits a systematic dependence on the schedule. Slower late-time decays expand the parameter regions where AQA outperforms TA, with the effect most pronounced at larger e. In contrast, faster decays yield comparatively larger regions where TA remains dominant. These observations are consistent with the view that, at fixed total runtime, emphasizing late-time evolution enhances AQA’s ability to resolve and traverse competing low-energy structures.
V Conclusion
This study investigated quantum and classical annealing dynamics for an open boundary spin-$1$ Ising chain with quadratic anisotropy. Across the $(J,D)$ plane, we observed a consistent pattern: quantum advantage emerges when two conditions are jointly satisfied, namely $(i)$ a sufficiently delocalized initial state (large $e$) and $(ii)$ schedules that allocate more time at small $g$. In these regimes, especially near order-change ridges where the dominant pattern shifts (all-0, all-$\pm 1$, antiferromagnetic), the dynamics benefit from stepwise rearrangements rather than a single collective switch. This distributes the evolution over multiple avoided crossings and can improve adiabatic transport relative to schedules that effectively force a single collective switch.
A second, schedule dependent effect concerns how the annealing time is distributed between the driver dominated and problem dominated parts of the evolution. For decays that naturally dwell near the final Hamiltonian (e.g., $g(t)\propto 1/\log(1{+}t)$) and at larger $e$, we observe systematically larger regions in which AQA outperforms TA, indicating that additional late-time budget is crucial for resolving competition between low energy patterns once the transverse field is weak. Linear ramps, by contrast, retain sizable TA dominant patches even at e=20e{=}20, underscoring that the terminal slowdown is an operational rather than purely cosmetic feature. Overall, TA performance follows the structure of classical barriers, while AQA shows additional sensitivity to schedule shape and transverse scale, consistent with coherent pathways that circumvent the need for collective flips (see App. A for supporting classical landscape diagnostics).
All comparisons use matched time budgets and identical $g(t)$ grids, so the observed advantage is a finite time effect under controlled conditions rather than an asymptotic separation. The present study nevertheless considers a single chain with open boundary condition, a fixed symmetry breaking field $h$, and one classical baseline. Future comparisons against additional classical heuristics (e.g., simulated annealing with optimized temperature schedules, population annealing, or cluster updates adapted to ternary variables) would strengthen the baseline. Likewise, scaling with system size and alternative graph topologies (rings, weak disorder, or sparse long-range couplings) will be important to assess robustness. Finally, schedule design itself is a practical lever: local adiabatic ramps, explicit end pauses, or counter-diabatic corrections could sharpen the quantum protocol without altering the problem Hamiltonian.
To the best of our knowledge, this is one of the first detailed investigations of ternary optimization within a quantum annealing framework using a spin-$1$ system with a quadratic anisotropy. We find evidence for a narrow but reproducible region in the $(J,D)$ plane where quantum systems outperform classical systems under matched finite time budgets. The advantage becomes most visible when the initial transverse scale $e$ is large and the schedule allocates sufficient time at small $g$, conditions that jointly promote coherent transport across sequences of avoided crossings near order-change boundaries, particularly in the easy-plane sector ($D>0$). Conversely, with insufficient driving or an aggressive terminal ramp, AQA can underperform even on classically easy instances.
These findings point to a pragmatic strategy for multi valued optimization: shape the problem’s local structure, particularly through the $D$ term, to open coherent pathways, and employ schedules that prioritize the endgame of the evolution. A more definitive characterization of the quantum classical performance boundary will require comprehensive spectral and landscape diagnostics, additional observables such as entanglement and participation ratios, and broader baselines and system sizes. Taken together, the present results and these next steps outline a concrete path toward establishing when and how spin-$1$ quantum annealing can provide a meaningful advantage in ternary optimization settings.
Acknowledgments
We thank Şeyda Leyla Bozan for suggesting useful and relevant references.
Appendix A Energy Landscape Structure
This appendix provides supporting evidence for the qualitative interpretation of landscape used in the main text. As a concrete proxy for how “funnel-like” versus fragmented the classical landscape is under local updates, we (i) quantify ruggedness through a basin decomposition induced by one-step moves and (ii) visualize the distribution of one-step local minima by plotting the classical energy $H(s)$ against a simple order parameter.
We define adjacency via one-step single site moves $s_{i}\mapsto s_{i}\pm 1$ (i.e. $1\leftrightarrow 0\leftrightarrow+1$). A configuration $m$ is a one-step local minimum if none of its one-step neighbors has strictly lower energy. Starting from any configuration $s$, we iteratively move to a strictly lower energy one-step neighbor (using a deterministic tie-break when multiple downhill moves exist) until reaching a fixed point $m$. This procedure partitions the full configuration space into basins of attraction $B(m)$, indexed by their corresponding local minima $m$.
We report two summary statistics:
  • Number of basins: the number of distinct one-step local minima.
  • Largest basin fraction: $\max_{m}|B(m)|/3^{N}$, the fraction of all configurations that drain into the largest basin.
These quantities are computed by enumeration to provide a concrete, interpretable proxy for landscape ruggedness.
Figure 2: Number of basins versus $D$ along an antiferromagnetic cut (J=2J=-2). Stepwise decreases reflect qualitative reorganizations of the set of minima as the anisotropy is varied.
Figure 3: Dominance of the largest basin versus $D$ along the cut (J=2J=-2). Growth of this fraction signals an increasingly funnel-like landscape; values near unity indicate that almost all configurations drain to a single minimum.
Figures 2 and 3 quantify how the classical landscape reorganizes with $D$ along a representative antiferromagnetic cut. For strongly negative $D$, many distinct local minima coexist and no single basin dominates, consistent with a rugged landscape. As $D$ increases, the basin count drops in discrete steps while the largest basin fraction rises, indicating a progressive shift toward a more funnel-like structure. For sufficiently large positive $D$, the largest basin fraction approaches unity and the basin count collapses to one, consistent with a classically simple limit.
To characterize these landscapes, we introduce the fraction of sites occupying the $\pm 1$ levels for a given classical configuration $s$,
$f(s)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{|s_{i}|=1},$ (20)
where $\mathbf{1}_{|s_{i}|=1}=1$ if $s_{i}=\pm 1$ and 0 otherwise. Thus $f=0$ for the all-0 configuration and $f=1$ for configurations with no 0-spins.
Figure 4: Easy-plane energy landscape along an antiferromagnetic cut. Classical energy $H(s)$ versus the fraction $f(s)$ of spins in the $\pm 1$ levels for J=2J=-2 and $D=+2$. Light blue points show all configurations; the solid line indicates, for each $f$, the minimum energy over configurations with that $f$; star markers denote one-step local minima (basin roots). The local minima separate into distinct clusters at different $f$, reflecting competing orders and a multi-basin structure in this order-parameter coordinate.
Figure 5: Easy-axis energy landscape along an antiferromagnetic cut. Classical energy $H(s)$ versus fraction $f(s)$ for J=2J=-2 and D=2D=-2, with one-step local minima marked by stars. In contrast to the easy-plane case, low energy states and local minima concentrate near $f=1$ and are less clearly separated in $f$, indicating a comparatively less partitioned structure in this coordinate.
Figures 4 and 5 visualize the distribution of one-step local minima in a physically meaningful coordinate. In the easy-plane case ($D=+2$), local minima (stars) occur in multiple well separated clusters at different $f$, consistent with competing orders involving both 0-rich and $\pm 1$-rich configurations. In the easy-axis case (D=2D=-2), the low energy configurations and minima are concentrated near $f=1$ and are less clearly separated in $f$, consistent with the 0-level being energetically disfavored and the low energy structure being dominated by $\pm 1$ occupations.

### v2

@@
 Quantum annealing offers a promising strategy for solving complex optimization problems by encoding the solution into the ground state of a problem Hamiltonian.
-While most implementations rely on spin-$1/2$ systems, we explore the performance of quantum annealing on a spin-$1$ system where the problem Hamiltonian includes a single ion anisotropy term of the form $D\sum(S^{z})^{2}$.
-Our results reveal that for a suitable range of the anisotropy strength $D$, the spin-$1$ annealer reaches the ground state with higher fidelity.
+While most implementations rely on spin-$1/2$ systems, we explore the performance of quantum annealing on a spin-$1$ system where the problem Hamiltonian includes a single-ion anisotropy term of the form $D\sum(S^{z})^{2}$ and compare it with trit based simulated annealing.
+Our results reveal that for a suitable range of the anisotropy strength $D$, the spin-$1$ annealer reaches the ground state with higher success probability.
 We attribute this performance to the presence of the intermediate spin level and the tunable anisotropy, which together enable the algorithm to traverse the energy landscape through smaller, incremental steps instead of a single large spin flip.
 This mechanism effectively lowers barriers in the configuration space and stabilizes the evolution.
-These findings suggest that higher spin annealers offer intrinsic advantages for robust and flexible quantum optimization, especially for problems naturally formulated with ternary decision variables.
+These findings suggest that higher-spin annealers can provide a finite-time performance benefit relative to the trit annealing baseline studied here, indicating a potentially robust and flexible approach to quantum optimization for problems naturally formulated with ternary decision variables.
 I Introduction
-Combinatorial optimization arises across physics, computer science, and engineering, with applications from materials and circuit design to logistics, portfolio optimization, and scheduling Harrison2022max3cutcoherent; Weggemans2022solvingcorrelation.
-Many such problems can be cast as minimizing a cost function over discrete variables, naturally mapping to the ground state of an effective Ising-type Hamiltonian with local fields and pairwise couplings Finnila1994; Kadowaki1998; Farhi2001; Das2008; Lucas2014.
+Combinatorial optimization arises across physics, computer science, and engineering, with applications from materials and circuit design to logistics, portfolio optimization, and scheduling Harrison et al.
+(2022); Weggemans et al.
+(2022).
+Many such problems can be cast as minimizing a cost function over discrete variables, naturally mapping to the ground state of an effective Ising-type Hamiltonian with local fields and pairwise couplings Finnila et al.
+(1994); Kadowaki and Nishimori (1998); Childs et al.
+(2001); Das and Chakrabarti (2008); Lucas (2014).
 Candidate solutions correspond to spin configurations, while optimal or near-optimal solutions lie at low energies.
 Since the configuration space grows exponentially with system size, classical methods can struggle on rugged, frustrated landscapes with many nearly degenerate local minima.
-Quantum annealing (QA) is a heuristic approach to such optimization problems that leverages quantum fluctuations Kadowaki1998; Farhi2001; Albash2018Review.
+Quantum annealing (QA) is a heuristic approach to such optimization problems that leverages quantum fluctuations Kadowaki and Nishimori (1998); Childs et al.
+(2001); Albash and Lidar (2018).
 The protocol initializes the system in the ground state of a simple driver Hamiltonian—typically a transverse field term that does not commute with the problem Hamiltonian—and then interpolates toward the problem Hamiltonian under a time-dependent schedule.
 In the coherent, adiabatic limit, the system follows the instantaneous ground state provided the evolution is sufficiently slow.
-In practice, the runtime is constrained by the minimum spectral gap encountered along the annealing path Farhi2001; Albash2018Review; Jansen2006BoundsFT; small gaps therefore require long schedules and increase susceptibility to nonadiabatic excitations, decoherence, and control errors Bottarelli2025symmetrycd.
-Most theoretical and experimental QA platforms use effective spin-$1/2$ degrees of freedom (qubits), which directly encode binary decision variables Das2008; Albash2018Review; Hauke2020; Johnson2011; DenchevPRX2016; BoixoNC2016.
+In practice, the runtime is constrained by the minimum spectral gap encountered along the annealing path Childs et al.
+(2001); Albash and Lidar (2018); Jansen et al.
+(2007); small gaps therefore require long schedules and increase susceptibility to nonadiabatic excitations, decoherence, and control errors Bottarelli et al.
+(2025).
+Most theoretical and experimental QA platforms use effective spin-$1/2$ degrees of freedom (qubits), which directly encode binary decision variables Das and Chakrabarti (2008); Albash and Lidar (2018); Hauke et al.
+(2020); Johnson et al.
+(2011); Denchev et al.
+(2016); Boixo et al.
+(2016).
 However, many practically relevant optimization problems are intrinsically multivalued, motivating approaches that go beyond binary encoding.
 Examples include ternary decisions (e.g., buy/hold/sell) and formulations with variables taking values in sets such as $\{-1,0,+1\}$.
-On qubit based architectures, representing multivalued logic typically requires embedding each variable into multiple qubits together with penalty terms enforcing logical constraints Weggemans2022solvingcorrelation.
+On qubit based architectures, representing multivalued logic typically requires embedding each variable into multiple qubits together with penalty terms enforcing logical constraints Weggemans et al.
+(2022).
 These embeddings increase hardware overhead, reduce effective connectivity, and complicate the energy landscape, potentially shrinking relevant gaps and amplifying sensitivity to noise and control imperfections.
@@
 Spin-11 systems, in particular, have three eigenstates of $S^{z}$ labeled by $m_{s}=-1,0,+1$, enabling a direct representation of ternary variables without elaborate embeddings.
-This has motivated interest in multilevel quantum optimization and in understanding how extra levels affect adiabatic protocols Amin2013Qudits; WangHuSandersKais2020; Weggemans2022solvingcorrelation; Bottarelli2025symmetrycd.
+This has motivated interest in multilevel quantum optimization and in understanding how extra levels affect adiabatic protocols Amin et al.
+(2013); Wang et al.
+(2020); Weggemans et al.
+(2022); Bottarelli et al.
+(2025).
 Beyond compact encoding, spin-$1$ models also introduce control parameters absent in spin-$1/2$ systems.
-A key example is the single ion anisotropy term $D\sum_{i}(S_{i}^{z})^{2}$, familiar from spin-$1$ magnetism such as the Blume–Capel model Blume1966; Capel1966, which tunes the relative energetics of the $m_{s}=0$ and $m_{s}=\pm 1$ states.
+A key example is the single-ion anisotropy term $D\sum_{i}(S_{i}^{z})^{2}$, familiar from spin-$1$ magnetism such as the Blume–Capel model Blume (1966); Capel (1966), which tunes the relative energetics of the $m_{s}=0$ and $m_{s}=\pm 1$ states.
 Since $(S^{z})^{2}$ is proportional to the identity for spin-$1/2$, this anisotropy provides a genuine additional control knob for spin-$1$ annealers and gives rise to richer behavior along the annealing path.
@@
   • $h$ is the longitudinal field, representing an external field that creates an energy preference for spin alignment in the $z$ direction,
-  • $D$ is the single ion anisotropy parameter,
+  • $D$ is the single-ion anisotropy parameter,
   • $g(t)$ is the annealing schedule controlling the strength of the transverse field.
@@
 $1-P_{gs}\lesssim\max_{t}\frac{\big|\langle\psi_{1}(t)|\dot{H}(t)|\psi_{0}(t)\rangle\big|^{2}}{\Delta(t)^{4}},$ (7)
-with gap $\Delta(t)=E_{1}(t)-E_{0}(t)$. Near an avoided crossing with minimum gap $\Delta_{\min}$, a two level linearization yields the Landau–Zener formula Landau
+with gap $\Delta(t)=E_{1}(t)-E_{0}(t)$. Near an avoided crossing with minimum gap $\Delta_{\min}$, a two level linearization yields the Landau–Zener formula Zener (1932)
 $P_{\rm exc}\approx\exp\left(-\frac{\pi\Delta_{\min}^{2}}{2\hbar|\alpha|}\right),$ (8)
 where $\alpha$ is the diabatic slope difference.
-The corresponding classical configuration space is
+We use the term trit for a classical three-valued variable $s\in\{-1,0,+1\}$. The corresponding classical configuration space is
 $\mathcal{S}=\{-1,0,+1\}^{N},$ (9)
@@
 $H(s)$ $=-J\sum_{\langle i,j\rangle}s_{i}s_{j}-h\sum_{i=1}^{N}s_{i}+D\sum_{i=1}^{N}s_{i}^{2}.$ (10)
-At temperature $T>0$, a single-site Metropolis update Metropolis1953 proceeds as follows:
+At temperature $T>0$, a single-site Metropolis update Metropolis et al. (1953) proceeds as follows:
   1. 1.
@@
 The transverse driver amplitude is scheduled as
-$g(t)=\frac{e}{f(t)},\qquad e\in\{1,5,10,20\},$ (16)
+$g(t)=\frac{c}{f(t)},\qquad c\in\{1,5,10,20\},$ (16)
 with four representative decay profiles
@@
 TA is defined on the same set of spin configurations by taking the classical cost function to be the diagonal of the quantum problem Hamiltonian, so that classical and quantum instances are exactly aligned at the level of their energy landscapes.
-One sweep comprises $N$ single site updates, in which each spin is updated once on average according to the local Boltzmann weights at the current effective temperature.
+One sweep comprises $N$ single-site updates, in which each spin is updated once on average according to the local Boltzmann weights at the current effective temperature.
 The inverse temperature is reduced following the same schedule shape $g(t)$ used in the quantum protocol, now applied over the course of $S$ sweeps.
-For each disorder instance (coupling realization) we perform $R$ independent TA runs starting from random initial spin configurations drawn uniformly from the $3^{N}$ classical states.
+For each instance, we perform $R$ independent TA runs starting from random initial spin configurations drawn uniformly from the $3^{N}$ classical states.
 At the end of each run $(r)$ we record the terminal energy $E_{\mathrm{final}}^{(r)}$.
@@
 where $\mathbb{I}[\cdot]$ denotes the indicator function.
-The small numerical tolerance accounts for floating point round off while effectively requiring that the algorithm reaches the true ground state.
-By matching instance sets, schedule shapes, and problem Hamiltonians between AQA and TA, we can directly compare their success probabilities.
+The small numerical tolerance accounts for floating-point round-off while effectively requiring that the algorithm reaches the true ground state.
 IV Results
 Figure 1: AQA versus TA across the coupling–anisotropy plane.
-Each panel displays the difference $P_{\mathrm{AQA}}-P_{\mathrm{TA}}$ over the $(J,D)$ plane for a fixed transverse field scale $e$ and annealing schedule $g(t)$, at longitudinal field $h=0.2$.
-Columns (left to right) correspond to schedules with late-time decay $g(t)\propto 1/\log(1{+}t)$, $1/\sqrt{t}$, $1/t$, and $1/t^{2}$, while rows (top to bottom) correspond to $e=1,5,10,20$.
+Each panel displays the difference $P_{\mathrm{AQA}}-P_{\mathrm{TA}}$ over the $(J,D)$ plane for a fixed transverse field scale $c$ and annealing schedule $g(t)$, at longitudinal field $h=0.2$.
+Columns (left to right) correspond to schedules with late-time decay $g(t)\propto 1/\log(1{+}t)$, $1/\sqrt{t}$, $1/t$, and $1/t^{2}$, while rows (top to bottom) correspond to $c=1,5,10,20$.
 Red (blue) indicates parameter regions where AQA yields higher (lower) success probability than TA, and yellow markers indicate cases with $P_{\mathrm{AQA}}>0.9$.
-The dashed gray lines at $J=0$ and $D=0$ delineate the easy-plane ($D>0$) and easy-axis ($D<0$) sectors and assist in identifying symmetry-related structures.
+The $(J,D)$ plane is sampled on a uniform grid with spacing $\Delta J=\Delta D=0.2$ over $J,D\in[-5,5]$.
+For each grid point, AQA is run for total runtime $T=1000$, and TA is run with an aligned budget of $S=1000$ sweeps and $R=20$ independent restarts.
+The dashed gray lines at $J=0$ and $D=0$ separate the easy-plane ($D>0$) and easy-axis ($D<0$) sectors and assist in identifying symmetry-related structures.
 Figure 1 compares the performance of AQA and TA across the $(J,D)$ plane for all transverse field scales and schedules considered.
-The dominant trends are governed by the driver strength $e$, the anisotropy sector (easy-plane $D>0$ versus easy-axis $D<0$), and the late-time decay of the schedule $g(t)$.
-For weak driving ($e=1$, top row), TA is comparable to AQA over nearly the entire parameter range, as indicated by the predominance of blue regions across all schedules.
+Yellow markers highlight instances with $P_{\mathrm{AQA}}>0.9$; this threshold is used only to visually identify near-unit success and does not affect the underlying heatmaps.
+The dominant trends are governed by the driver strength $c$, the anisotropy sector (easy-plane $D>0$ versus easy-axis $D<0$), and the late-time decay of the schedule $g(t)$.
+We stress that throughout the paper the statement “AQA outperforms TA” is used in a strictly operational sense, i.e., $P_{\mathrm{AQA}}>P_{\mathrm{TA}}$ for the particular TA baseline considered here (local single-spin Metropolis updates for trits) under matched finite-time budgets, and should not be interpreted as a general separation from all classical annealing heuristics.
+For weak driving ($c=1$, top row), TA is comparable to AQA over nearly the entire parameter range, as indicated by the predominance of blue regions across all schedules.
 In this regime, the transverse field contribution is insufficient, within the fixed runtime, to reliably generate transitions that improve the final success probability.
 The sparse red features occur in narrow parameter bands near boundaries between distinct classical ordering tendencies, consistent with cases where near degeneracies permit limited benefit from coherent mixing.
-As the driver strength increases ($e=5$ and above), extended parameter regions emerge in which AQA outperforms TA.
-This advantage is most apparent in the easy-plane sector ($D>0$) at intermediate coupling magnitudes $|J|$, where red domains become substantial for $e=5$ and expand further for $e=10$ and $e=20$.
+As the driver strength increases ($c=5$ and above), extended parameter regions emerge in which AQA outperforms TA.
+This advantage is most apparent in the easy-plane sector ($D>0$) at intermediate coupling magnitudes $|J|$, where red domains become substantial for $c=5$ and expand further for $c=10$ and $c=20$.
 The growth of these AQA favorable regions indicates that sufficiently strong transverse driving can more effectively exploit low energy structure associated with competing classical patterns.
@@
 These boundary regions are frequently associated with reduced performance for both protocols, producing narrow corridors in which the relative advantage alternates and separating neighboring domains with more uniform behavior.
-At fixed e, performance exhibits a systematic dependence on the schedule.
-Slower late-time decays expand the parameter regions where AQA outperforms TA, with the effect most pronounced at larger e.
+At fixed $c$, performance exhibits a systematic dependence on the schedule.
+Slower late-time decays expand the parameter regions where AQA outperforms TA, with the effect most pronounced at larger $c$.
 In contrast, faster decays yield comparatively larger regions where TA remains dominant.
@@
 This study investigated quantum and classical annealing dynamics for an open boundary spin-$1$ Ising chain with quadratic anisotropy.
-Across the $(J,D)$ plane, we observed a consistent pattern: quantum advantage emerges when two conditions are jointly satisfied, namely $(i)$ a sufficiently delocalized initial state (large $e$) and $(ii)$ schedules that allocate more time at small $g$.
+Across the $(J,D)$ plane, we observed a consistent pattern: quantum advantage emerges when two conditions are jointly satisfied, namely $(i)$ a sufficiently delocalized initial state (large $c$) and $(ii)$ schedules that allocate more time at small $g$.
 In these regimes, especially near order-change ridges where the dominant pattern shifts (all-0, all-$\pm 1$, antiferromagnetic), the dynamics benefit from stepwise rearrangements rather than a single collective switch.
@@
 A second, schedule dependent effect concerns how the annealing time is distributed between the driver dominated and problem dominated parts of the evolution.
-For decays that naturally dwell near the final Hamiltonian (e.g., $g(t)\propto 1/\log(1{+}t)$) and at larger $e$, we observe systematically larger regions in which AQA outperforms TA, indicating that additional late-time budget is crucial for resolving competition between low energy patterns once the transverse field is weak.
-Linear ramps, by contrast, retain sizable TA dominant patches even at e=20e{=}20, underscoring that the terminal slowdown is an operational rather than purely cosmetic feature.
+For decays that naturally dwell near the final Hamiltonian (e.g., $g(t)\propto 1/\log(1{+}t)$) and at larger $c$, we observe systematically larger regions in which AQA outperforms TA, indicating that additional late-time budget is crucial for resolving competition between low energy patterns once the transverse field is weak.
+Linear ramps, by contrast, retain sizable TA dominant patches even at c=20c{=}20, underscoring that the terminal slowdown is an operational rather than purely cosmetic feature.
 Overall, TA performance follows the structure of classical barriers, while AQA shows additional sensitivity to schedule shape and transverse scale, consistent with coherent pathways that circumvent the need for collective flips (see App.
 A for supporting classical landscape diagnostics).
-All comparisons use matched time budgets and identical $g(t)$ grids, so the observed advantage is a finite time effect under controlled conditions rather than an asymptotic separation.
+All comparisons use matched time budgets and identical $g(t)$ grids, so the observed advantage is a finite-time effect under controlled conditions rather than an asymptotic separation.
 The present study nevertheless considers a single chain with open boundary condition, a fixed symmetry breaking field $h$, and one classical baseline.
@@
 To the best of our knowledge, this is one of the first detailed investigations of ternary optimization within a quantum annealing framework using a spin-$1$ system with a quadratic anisotropy.
-We find evidence for a narrow but reproducible region in the $(J,D)$ plane where quantum systems outperform classical systems under matched finite time budgets.
-The advantage becomes most visible when the initial transverse scale $e$ is large and the schedule allocates sufficient time at small $g$, conditions that jointly promote coherent transport across sequences of avoided crossings near order-change boundaries, particularly in the easy-plane sector ($D>0$).
+We find evidence for a narrow but reproducible region in the $(J,D)$ plane where quantum systems outperform classical systems under matched finite-time budgets.
+The advantage becomes most visible when the initial transverse scale $c$ is large and the schedule allocates sufficient time at small $g$, conditions that jointly promote coherent transport across sequences of avoided crossings near order-change boundaries, particularly in the easy-plane sector ($D>0$).
 Conversely, with insufficient driving or an aggressive terminal ramp, AQA can underperform even on classically easy instances.